# CALT-TH 2023-052

### 15 Dec 2023 v1

@@ A.1 The Linear Program and its Dual
To construct a linear program, we need to specify a set of variables, a set of linear constraints, and the extremization goal. In this subsection, we will assume our goal is to compute ${\texttt{S}}({\underline{I}})$, but the generalization to multipartite regions should be clear.
@@ A.2 Vertex Cuts
For every multi-cycle flow ${\mathfrak{C}}({\underline{I}})$ maximizing the flux through boundary vertices ${\underline{I}}$, there exists a vertex cut ${\mathcal{V}}_{{\underline{I}}}$ whose magnitude is precisely the max flow ${\texttt{S}}({\underline{I}})$, which we shall construct as follows. From ${\mathfrak{C}}({\underline{I}})$, we can construct a reduced multi-cycle flow ${\mathfrak{C}}^{r}({\underline{I}})$ by removing any simple cycle flow ${\mathcal{C}}\in{\mathfrak{C}}({\underline{I}})$ that does not contain a vertex in both ${\underline{I}}$ and ${\underline{I}}^{c}$. Notice that ${\mathfrak{C}}^{r}({\underline{I}})$ still maximizes the flow out of the ${\underline{I}}$ vertices; it just does not contain any extraneous simple cycle flows. Using ${\mathfrak{C}}^{r}({\underline{I}})$, we can construct the residual graph ${\text{Res}}({\mathcal{N}},{\mathfrak{C}}^{r}({\underline{I}}))$, as defined in (2.22).
In the residual graph ${\text{Res}}({\mathcal{N}},{\mathfrak{C}}^{r}({\underline{I}}))$, the vertices in ${\underline{I}}$ are disconnected from those in ${\underline{I}}^{c}$. Otherwise, either the network ${\mathcal{N}}$ is unbalanced, or the flux through the multi-cycle flow ${\mathfrak{C}}^{r}({\underline{I}})$ could be increased, which is not possible as it is a max cycle flow through ${\underline{I}}$. Consequently, we associate to ${\mathfrak{C}}({\underline{I}})$ the vertex cut ${\mathcal{V}}_{{\mathfrak{C}}({\underline{I}})}$ consisting of ${\underline{I}}$ and any other vertices (both internal and boundary) still connected to a vertex in ${\underline{I}}$ in the residual graph ${\text{Res}}({\mathcal{N}},{\mathfrak{C}}^{r}({\underline{I}}))$. As for the magnitude $|{\mathcal{V}}_{{\mathfrak{C}}({\underline{I}})}|$ of the cut, note that by (A.7), it cannot be simply the sum of the capacities of all edges leaving the cut, as the standard “min cut” perspective would suggest. [^15] Consider the 4-party GHZ state in Figure 5. If we choose our vertex cut to consist of vertices $\{A,C\}$, then by (A.7) $|{\mathcal{V}}(\{A,C\})|=1$. However, if we just summed the capacity of the edges which leave the vertex cut, the flow would instead be $2$. Instead, by (A.7), we see that
@@
Notice that this value coincides with the magnitude of the max flow $f_{{\mathfrak{C}}({\underline{I}})}({\underline{I}},{\underline{I}}^{c})$, since it was built from the residual network ${\text{Res}}({\mathcal{N}},{\mathfrak{C}}^{r}({\underline{I}}))$. Indeed, if $|{\mathcal{V}}_{{\mathfrak{C}}({\underline{I}})}|$ were larger than $f_{{\mathfrak{C}}({\underline{I}})}({\underline{I}},{\underline{I}}^{c})$, then the magnitude of the original multi-cycle flow ${\mathfrak{C}}({\underline{I}})$ could have been increased, which is impossible since it is a max flow arrangement. By strong duality, the magnitude of the max flow, which as we just argued is $|{\mathcal{V}}_{{\mathfrak{C}}({\underline{I}})}|$, is precisely the minimum value for (A.4).
@@ A.3 Alternative Proofs of SA
$S({\underline{I}})+S({\underline{J}})=|{\mathcal{V}}_{\underline{I}}|+|{\mathcal{V}}_{\underline{J}}|\geq|{\mathcal{V}}({\underline{I}})\cup{\mathcal{V}}({\underline{J}})|\geq|{\mathcal{V}}({\underline{I}}{\underline{J}})|={\texttt{S}}({\underline{I}}{\underline{J}}),$ (A.12)
@@
Of course, we are also interested to prove SSA from the dual perspective. At the moment, it is not clear which of the two perspectives, or perhaps a combination of the two, is the best way to prove SSA. We leave such explorations to future work.

### 12 Oct 2024 v3

@@ A.1 The Linear Program and its Dual
To construct a linear program, we need to specify a set of variables, a set of linear constraints, and the extremization goal. In this subsection, we will assume our goal is to compute ${\texttt{S}}({\underline{I}})$, but the generalization to multipartite regions ${\underline{I}}{\underline{J}}$ for any ${\underline{J}}\subset{\underline{I}}^{c}$ should be clear.
@@ A.2 Vertex Cuts
For every multi-cycle flow ${\mathfrak{C}}({\underline{I}})$ maximizing the flux through boundary vertices ${\underline{I}}$, there exists a vertex cut ${\mathcal{V}}_{{\underline{I}}}$ whose magnitude is precisely the max flow ${\texttt{S}}({\underline{I}})$, which we shall construct as follows. From ${\mathfrak{C}}({\underline{I}})$, we can construct a reduced multi-cycle flow ${\mathfrak{C}}^{r}({\underline{I}})$ by removing any simple cycle flow ${\mathcal{C}}\in{\mathfrak{C}}({\underline{I}})$ that does not contain a vertex in both ${\underline{I}}$ and ${\underline{I}}^{c}$. Notice that ${\mathfrak{C}}^{r}({\underline{I}})$ still maximizes the flow out of the ${\underline{I}}$ vertices; it just does not contain any extraneous simple cycle flows. Using ${\mathfrak{C}}^{r}({\underline{I}})$, we can construct the residual network ${\text{Res}}({\mathcal{N}},{\mathfrak{C}}^{r}({\underline{I}}))$, as defined in (2.22).
In the residual network ${\text{Res}}({\mathcal{N}},{\mathfrak{C}}^{r}({\underline{I}}))$, the vertices in ${\underline{I}}$ are disconnected from those in ${\underline{I}}^{c}$. Otherwise, either the network ${\mathcal{N}}$ is unbalanced, or the flux through the multi-cycle flow ${\mathfrak{C}}^{r}({\underline{I}})$ could be increased, which is not possible as it is a max cycle flow through ${\underline{I}}$. Consequently, we associate to ${\mathfrak{C}}({\underline{I}})$ the vertex cut ${\mathcal{V}}_{{\mathfrak{C}}({\underline{I}})}$ consisting of ${\underline{I}}$ and any other vertices (both internal and boundary) still connected to a vertex in ${\underline{I}}$ in the residual network ${\text{Res}}({\mathcal{N}},{\mathfrak{C}}^{r}({\underline{I}}))$. As for the magnitude $|{\mathcal{V}}_{{\mathfrak{C}}({\underline{I}})}|$ of the cut, note that by (A.7), it cannot be simply the sum of the capacities of all edges leaving the cut, as the standard “min cut” perspective would suggest. [^19] Consider the 4-party GHZ state in Figure 6. If we choose our vertex cut to consist of vertices $\{A,C\}$, then by (A.7) $|{\mathcal{V}}(\{A,C\})|=1$. However, if we just summed the capacity of the edges which leave the vertex cut, the flow would instead be $2$. Instead, by (A.7), we see that
@@
Notice that this value coincides with the magnitude of the max flow $f_{{\mathfrak{C}}({\underline{I}})}({\underline{I}},{\underline{I}}^{c})$, since it was built from the residual network ${\text{Res}}({\mathcal{N}},{\mathfrak{C}}^{r}({\underline{I}}))$. Indeed, if $|{\mathcal{V}}_{{\mathfrak{C}}({\underline{I}})}|$ were larger than $f_{{\mathfrak{C}}({\underline{I}})}({\underline{I}},{\underline{I}}^{c})$, then the magnitude of the original multi-cycle flow ${\mathfrak{C}}({\underline{I}})$ could have been increased, which is impossible since it is a max flow arrangement. This proves ${\texttt{S}}({\underline{I}})$ is the magnitude of one such vertex cut containing ${\underline{I}}$, and we denote this vertex cut ${\mathcal{V}}_{\underline{I}}$.
To see that ${\texttt{S}}({\underline{I}})$ is the minimum magnitude of all cuts containing only the boundary vertices ${\underline{I}}$ (but may contain additional internal vertices), suppose ${\mathcal{V}}^{\prime}_{\underline{I}}$ is another such vertex cut. If the vertex set ${\mathcal{V}}^{\prime}_{\underline{I}}$ is disconnected from its complement $V\setminus{\mathcal{V}}^{\prime}_{\underline{I}}$ in ${\text{Res}}({\mathcal{N}},{\mathfrak{C}}^{r}({\underline{I}}))$, then ${\mathfrak{C}}^{r}({\underline{I}})$ is one such flow that we are maximizing over in (A.7) when determining $|{\mathcal{V}}^{\prime}_{\underline{I}}|$. Otherwise, we have to append additional simple cycles to ${\mathfrak{C}}^{r}({\underline{I}})$, obtaining a resultant multi-cycle flow ${\mathfrak{C}}^{\prime}({\underline{I}})$, to ensure ${\mathcal{V}}^{\prime}_{\underline{I}}$ is disconnected from its complement in the residual graph ${\text{Res}}({\mathcal{N}},{\mathfrak{C}}^{\prime}({\underline{I}}))$. In this case, ${\mathfrak{C}}^{\prime}({\underline{I}})$ is one such flow that we are maximizing over in (A.7) when determining $|{\mathcal{V}}^{\prime}_{\underline{I}}|$, and note that by construction
$f_{{\mathfrak{C}}^{\prime}({\underline{I}})}({\mathcal{V}}^{\prime}_{\underline{I}},V\setminus{\mathcal{V}}^{\prime}_{\underline{I}})\geq f_{{\mathfrak{C}}({\underline{I}})}({\mathcal{V}}_{\underline{I}},V\setminus{\mathcal{V}}_{\underline{I}}).$ (A.9)
Thus, we see that $|{\mathcal{V}}^{\prime}_{\underline{I}}|\geq|{\mathcal{V}}_{\underline{I}}|$, proving the claim.
@@ A.3 Alternative Proofs of SA
${\texttt{S}}({\underline{I}})+{\texttt{S}}({\underline{J}})=|{\mathcal{V}}({\underline{I}})|+|{\mathcal{V}}({\underline{J}})|\geq|{\mathcal{V}}({\underline{I}})\cup{\mathcal{V}}({\underline{J}})|\geq|{\mathcal{V}}({\underline{I}}{\underline{J}})|={\texttt{S}}({\underline{I}}{\underline{J}}),$ (A.13)
@@
Of course, we are also interested to prove SSA using the dual program. As we proved that SSA holds only for nesting balanced networks, this would involve understanding how to capture the notion of nesting cycle flows from the dual perspective. We leave such explorations to future work.